# Rethinking Multiple Instance Learning for Whole Slide Image Classification: A Good Instance Classifier is All You Need

### v1

@@ 4.1 Datasets
4.1.1 Simulated CIFAR-MIL Dataset
@@
The CIFAR-10 dataset consists of 60,000 32$\times$32-pixel color images divided into 10 categories (airplane, automobile, bird, cat, deer, dog, frog, horse, ship, truck), with each category containing 6,000 images. Out of these images, 50,000 are used for training and 10,000 for testing. To simulate pathological Whole Slide Images (WSIs), we combined a set of random images from each category of the CIFAR-10 dataset to construct the WSI. Specifically, we treated each image from each category of the CIFAR-10 dataset as an instance, and only all instances of the ”truck” category were labeled as positive while the other instances were labeled as negative (the truck category is chosen at random). Then, we randomly selected a positive bag consisting of a set of $a$ positive instances and $100-a$ negative instances (without repetition) from all instances, with a positive ratio of $\frac{a}{100}$. Similarly, we randomly selected a negative bag consisting of 100 negative instances (without repetition). We repeated this process until all positive or negative instances in the CIFAR-10 dataset were used up. By adjusting the value of a, we constructed 5 subsets of the CIFAR-MIL dataset with positive ratios of 5%, 10%, 20%, 50%, and 70%, respectively.
4.1.2 Camelyon16 Public Dataset
@@
4.1.3 TCGA Lung Cancer Dataset
@@
4.1.4 Cervical Cancer Dataset
@@
4.2 Evaluation Metrics and Comparison Methods
@@ 4.3 Implementation Details
4.4 Results on the Synthetic Dataset CIFAR-MIL
We constructed the synthetic WSI datasets CIFAR-MIL with varying positive instance ratios to investigate the instance and bag classification performance of each method at different positive instance ratios. The results are shown in Table 1 and Table 2. It can be seen that INS achieved the best performance in both instance and bag classification tasks at all positive instance ratios. Most methods cannot work well in low positive instance ratios of 5% and 10%, for which the AUC of INS exceeds 0.94. At positive instance ratios above 20%, the performance of INS is comparable to that of fully supervised methods.
@@ 4.4 Results on the Synthetic Dataset CIFAR-MIL
Table 1: Instance classification results on the CIFAR-MIL Dataset.
@@
Table 2: Bag classification results on the CIFAR-MIL Dataset.
@@
Table 3: Results on the Camelyon 16 Dataset.
Instance-level Bag-level
Methods AUC Accuracy AUC Accuracy
@@
Table 4: Results on the TCGA-Lung Cancer Dataset.
Bag-level
Methods AUC Accuracy
@@
Table 5: Results on the Cervical Cancer Dataset.
Clinical Tasks (Bag level) Lymph node metastasis Survival Prognosis KI-67 Prediction
Methods AUC Accuracy AUC Accuracy AUC Accuracy
ABMIL (18’ICML) 0.7716 0.7297 0.7439 0.7125 0.6963 0.6500
Loss-ABMIL (20’AAAI) 0.7835 0.7568 0.7505 0.7250 0.6987 0.6625
Chi-MIL (20’MICCAI) 0.7783 0.7432 0.7349 0.7000 0.6933 0.6625
DSMIL (21’CVPR) 0.8022 0.7838 0.7622 0.7375 0.7124 0.6875
TransMIL (21’NeurIPS) 0.8126 0.7838 0.7716 0.7375 0.7252 0.6875
CLAM (21’Nat. Biomed. Eng.) 0.8138 0.7838 0.7755 0.7375 0.7297 0.7000
DTFD-MIL (22’CVPR) 0.8108 0.7703 0.7650 0.7375 0.7267 0.6875
DGMIL (22’MICCAI) 0.8056 0.7568 0.7515 0.7250 0.7115 0.6875
WENO (22’NeurIPS) 0.8222 0.7703 0.7775 0.7500 0.7323 0.7000
TPMIL (23’MIDL) 0.8213 0.7703 0.7728 0.7500 0.7318 0.7000
@@
4.5 Results on Real-World Datasets
4.5.1 Camelyon 16 Dataset
Table 3 shows the classification performance of INS and other methods on the Camelyon 16 Dataset. The low positive ratio of Camelyon 16 Dataset (about 10%-20%) makes the classification task quite difficult. INS outperforms all the compared methods with a large margin, exceeding the second-best method by 2.1% and 3.8% in AUC for instance and bag classification, respectively. Figure 4 shows typical visualization results on this dataset. It can be seen that INS accurately localizes almost all positive instances in the positive bags.
4.5.2 TCGA-Lung Cancer Dataset
Table 4 shows the results on the TCGA-Lung Cancer Dataset. Unlike the Camelyon 16 Dataset, this dataset has a high positive instance ratio (more than 80%), so the performance of all methods is fairly good, while INS still achieves the best performance.
4.5.3 Cervical Cancer Dataset
Table 5 shows the experimental results of three tasks on the in-house Cervical Cancer Dataset, including lymph node metastasis prediction from primary lesion, patient prognosis, and immunohistochemical marker KI-67 prediction from H&E stained WSIs. Unlike the previous two real-world datasets, these three tasks are very tough even for human doctors. It can be seen that INS significantly outperforms other methods in all three tasks, demonstrating the strong performance of INS.
@@ 4.5.3 Cervical Cancer Dataset
Figure 5: Heatmap visualization of INS on the lymph node metastasis task, where we identify the ”Micropapillae” pathological pattern that indicates positive metastasis from HE slides.
4.6 Interpretability Study of the Lymph Node Metastasis Task
As can be seen from Table 5, INS can predict the preoperative lymph node status of cervical cancer with high performance using HE-stained slides of the primary tumor. However, this does not offer pathologists additional interpretable pathological patterns. While cancer diagnosis is possible through visual inspection of HE-stained pathological slides, the same cannot be said for accurately determining lymph node metastasis status from the primary lesion. Given INS’s powerful instance classification ability, we hope to find explicit pathological patterns that can suggest the lymph node metastasis status and perform interpretable analysis as well as new knowledge discovery. Specifically, we used INS to predict the probability of each instance being positive within the positive bags, and visualized the top 0.1% instances with the highest and lowest probabilities separately. These instances with the highest predicted positive probabilities are likely to contain pathological patterns that identify high-risk lymph node metastasis, as shown in Figure 5.
As can be seen, in images suggesting lymph node metastasis, structures resembling ”micropapillae” are more prevalent, indicating a high-risk pathological pattern. Micropapillary structures are characterized by small clusters of infiltrating cancer cells forming hollow or mulberry-like nests without a central fibrovascular axis, surrounded by blank lacunae or lacunae between interstitial components. Conversely, negative lymph node images more commonly exhibit a ”sheet-like” pattern, in which cells form tightly connected nests of varying sizes within the tumor interstitium, with fissures or lacunae rarely observed. This conclusion highlights the interpretability of using INS to assess lymph node metastasis and its significant guiding implications for future clinical research.
5 Ablation Study and Further Analysis
5.1 Ablation Study on Key Components
@@
Table 6: Results of ablation study on the Camelyon 16 Dataset.
@@ 5.2 Further Analysis
6 Conclusion

### v2

@@ IV-A Datasets
IV-A1 Simulated CIFAR-MIL Dataset
@@
The CIFAR-10 dataset consists of 60,000 32$\times$32-pixel color images divided into 10 categories (airplane, automobile, bird, cat, deer, dog, frog, horse, ship, truck), with each category containing 6,000 images. Out of these images, 50,000 are used for training and 10,000 for testing. To simulate pathological Whole Slide Images (WSIs), we combined a set of random images from each category of the CIFAR-10 dataset to construct the WSI. Specifically, we treated each image from each category of the CIFAR-10 dataset as an instance, and only all instances of the "truck" category were labeled as positive while the other instances were labeled as negative (the truck category is chosen at random). Then, we randomly selected a positive bag consisting of a set of $a$ positive instances and $100-a$ negative instances (without repetition) from all instances, with a positive ratio of $\frac{a}{100}$. Similarly, we randomly selected a negative bag consisting of 100 negative instances (without repetition). We repeated this process until all positive or negative instances in the CIFAR-10 dataset were used up. By adjusting the value of a, we constructed 5 subsets of the CIFAR-MIL dataset with positive ratios of 5%, 10%, 20%, 50%, and 70%, respectively.
IV-A2 Camelyon16 Public Dataset
@@
IV-A3 TCGA Lung Cancer Dataset
@@
IV-A4 Cervical Cancer Dataset
@@
IV-B Evaluation Metrics and Comparison Methods
@@ IV-C Implementation Details
IV-D Results on the Synthetic Dataset CIFAR-MIL
We constructed the synthetic WSI datasets CIFAR-MIL with varying positive instance ratios to investigate the instance and bag classification performance of each method at different positive instance ratios. The results are shown in Table I and Table II. It can be seen that INS achieved the best performance in both instance and bag classification tasks at all positive instance ratios. Most methods cannot work well in low positive instance ratios of 5% and 10%, for which the AUC of INS exceeds 0.94. At positive instance ratios above 20%, the performance of INS is comparable to that of fully supervised methods.
@@ IV-D Results on the Synthetic Dataset CIFAR-MIL
TABLE I: Instance classification results on the CIFAR-MIL Dataset.
@@
TABLE II: Bag classification results on the CIFAR-MIL Dataset.
@@
TABLE III: Results on the Camelyon 16 Dataset.
[…]
@@
TABLE IV: Results on the TCGA-Lung Cancer Dataset.
[…]
@@
TABLE V: Results on the Cervical Cancer Dataset.
[…]
@@
IV-E Results on Real-World Datasets
IV-E1 Camelyon 16 Dataset
Table III shows the classification performance of INS and other methods on the Camelyon 16 Dataset. The low positive ratio of Camelyon 16 Dataset (about 10%-20%) makes the classification task quite difficult. INS outperforms all the compared methods with a large margin, exceeding the second-best method by 2.1% and 3.8% in AUC for instance and bag classification, respectively. Figure 4 shows typical visualization results on this dataset. It can be seen that INS accurately localizes almost all positive instances in the positive bags.
IV-E2 TCGA-Lung Cancer Dataset
Table IV shows the results on the TCGA-Lung Cancer Dataset. Unlike the Camelyon 16 Dataset, this dataset has a high positive instance ratio (more than 80%), so the performance of all methods is fairly good, while INS still achieves the best performance.
IV-E3 Cervical Cancer Dataset
Table V shows the experimental results of three tasks on the in-house Cervical Cancer Dataset, including lymph node metastasis prediction from primary lesion, patient prognosis, and immunohistochemical marker KI-67 prediction from H&E stained WSIs. Unlike the previous two real-world datasets, these three tasks are very tough even for human doctors. It can be seen that INS significantly outperforms other methods in all three tasks, demonstrating the strong performance of INS.
@@ IV-E3 Cervical Cancer Dataset
Figure 5: Heatmap visualization of INS on the lymph node metastasis task, where we identify the "Micropapillae" pathological pattern that indicates positive metastasis from HE slides.
IV-F Interpretability Study of the Lymph Node Metastasis Task
As can be seen from Table V, INS can predict the preoperative lymph node status of cervical cancer with high performance using HE-stained slides of the primary tumor. However, this does not offer pathologists additional interpretable pathological patterns. While cancer diagnosis is possible through visual inspection of HE-stained pathological slides, the same cannot be said for accurately determining lymph node metastasis status from the primary lesion. Given INS’s powerful instance classification ability, we hope to find explicit pathological patterns that can suggest the lymph node metastasis status and perform interpretable analysis as well as new knowledge discovery. Specifically, we used INS to predict the probability of each instance being positive within the positive bags, and visualized the top 0.1% instances with the highest and lowest probabilities separately. These instances with the highest predicted positive probabilities are likely to contain pathological patterns that identify high-risk lymph node metastasis, as shown in Figure 5.
As can be seen, in images suggesting lymph node metastasis, structures resembling "micropapillae" are more prevalent, indicating a high-risk pathological pattern. Micropapillary structures are characterized by small clusters of infiltrating cancer cells forming hollow or mulberry-like nests without a central fibrovascular axis, surrounded by blank lacunae or lacunae between interstitial components. Conversely, negative lymph node images more commonly exhibit a "sheet-like" pattern, in which cells form tightly connected nests of varying sizes within the tumor interstitium, with fissures or lacunae rarely observed. This conclusion highlights the interpretability of using INS to assess lymph node metastasis and its significant guiding implications for future clinical research.
V Ablation Study and Further Analysis
V-A Ablation Study on Key Components
@@
TABLE VI: Results of ablation study on the Camelyon 16 Dataset.
@@ V-B Further Analysis
VI Discussion
VI-A Purpose and Clinical Relevance of Instance vs. Bag Classification in IWSCL.
Purpose and Meaning of IWSCL: Instance-level Weakly Supervised Contrastive Learning (IWSCL) is designed to enhance the performance of MIL in the context of WSI analysis by improving the representation of instances (small patches derived from WSIs) in feature space. Traditional MIL approaches suffer from the challenge of accurately labeling each instance and identifying difficult positive instances. IWSCL addresses this by learning discriminative feature representations that better separate positive and negative instances, thereby facilitating more accurate pseudo-labeling and instance classification.
Definition of Instances and Their Importance: In our study, instances refer to the small patches into which a WSI is divided for analysis. Each instance represents a fraction of the larger slide, and its correct classification is crucial for understanding the slide’s overall pathology. The distinction between positive (indicative of disease) and negative (healthy) instances is foundational to accurate diagnosis at both the instance and slide levels.
Clinical Relevance of Bag vs. Instance Classification: Bag-Level Classification refers to the classification of the entire WSI (the "bag" in MIL terminology) as either positive or negative based on the aggregated information from its instances. Clinically, this is relevant for determining the presence or absence of disease in the slide as a whole, which is a crucial first step in diagnostic workflows. Instance-Level Classification involves classifying each patch (instance) within a WSI. Clinically, this is crucial for identifying specific areas of interest, such as tumor regions, and understanding the heterogeneity within a slide. It is particularly relevant for detailed analyses, such as grading the severity of disease or guiding targeted biopsies. While both levels of classification are clinically relevant, their importance varies with the diagnostic task. Bag-level classification is essential for initial disease presence screening, whereas instance-level classification plays a key role in detailed pathological assessments and understanding disease mechanisms.
To conclude, our work on IWSCL and the proposed MIL framework aims to improve both bag and instance-level classifications by addressing the limitations of existing methods. By enhancing the accuracy of instance classification, we indirectly improve bag-level classification outcomes, thus offering significant benefits for pathological diagnosis and analysis. We believe that our approach, by improving the granularity and accuracy of classifications, holds substantial promise for clinical applications, where precise localization and characterization of pathological features are critical.
VI-B Reliability of Pseudo-labels and the Instance Classifier
The accuracy of pseudo-labels is indeed critical in our approach, given their direct influence on the training of the instance classifier. We have devised a comprehensive strategy to ensure the reliability and accuracy of these pseudo-labels.
Prototype-based Pseudo Label Generation (PPLG): Our PPLG strategy is at the heart of ensuring the precision of pseudo-labels. By maintaining two dynamic prototypes representing negative and positive instances within the feature space, we can generate high-quality pseudo-labels for each instance based on their proximity to these prototypes. This method significantly mitigates the pseudo-label noise by leveraging the global context of instance distributions rather than individual, isolated predictions.
Contrastive Learning for Enhanced Feature Representation: Prior to pseudo-label generation, our framework employs contrastive learning to refine the feature representations of instances. This step is crucial for ensuring that positive and negative instances are well-separated in the feature space, thereby enhancing the subsequent pseudo-label accuracy. Improved feature representation directly contributes to the reliability of PPLG and, by extension, the precision of the instance classifier.
Joint Training Approach: The training process of the instance classifier is intricately linked with both IWSCL for feature representation enhancement and PPLG for pseudo-label generation. This interconnected training regime ensures a synergistic improvement across all components. As the instance classifier becomes more accurate, it further refines the quality of feature representations and pseudo-labels, creating a positive feedback loop that elevates the overall system performance.
Utilization of True Negative Instances: An innovative aspect of our methodology is the strategic use of true negative instances from negative bags. These instances provide a stable reference point for the model, helping to anchor the pseudo-label generation process and ensure that our classifier is calibrated against known standards, thereby improving its predictive accuracy.
Experimental results also directly validate the quality of the instance pseudo labels. As shown in Section V-B "Evaluation of Pseudo Labels", we further plotted the AUC curve of the pseudo labels for all instances in the training set of the CIFAR-MIL Dataset (with a 0.2 positive instance ratio) against the training epochs, as shown in Figure 6 B. It can be clearly seen that the quality of the pseudo labels continues to improve, providing more and more effective guidance to the instance classifier.
To conclude, our comprehensive approach, characterized by PPLG, contrastive learning for feature refinement, a synergistic training strategy, and the strategic use of true negatives, is meticulously designed to ensure the accuracy of pseudo-labels and the predictive reliability of the instance classifier. These mechanisms collectively address the challenges associated with pseudo-label noise and classifier training, thereby significantly enhancing the performance of our proposed framework in WSI classification tasks. Experimental results also directly validate the quality of the instance pseudo labels.
VI-C Novelty of the Proposed IWSCL
Innovation in WSI Classification: Our innovation primarily arises from integrating established components of contrastive learning tailored specifically for weakly supervised instance-level WSI classification. While mechanisms like loss terms and momentum updates are well-known in the field of contrastive learning, their application within an instance-level weakly supervised context for WSI analysis is novel. The unique challenges presented by WSIs, such as their large size, weak annotation, detailed content, and significant variance in pathological features, necessitate specialized solutions. We address these by adapting and optimizing contrastive learning concepts to the specific requirements of WSI classification, thus advancing their application in the field of WSI analysis.
Comparison with Existing Methods: As discussed in Section II-C, current methodologies typically commence with self-supervised learning on WSI patches to pretrain a feature extractor, followed by training using these features. Most utilize unsupervised contrastive learning to extract instance features but fail to specifically separate positive from negative instances. In contrast, our approach introduces IWSCL under the MIL framework for the first time, effectively distinguishing between negative and positive instance features, as illustrated in Figure 2.
Role in Joint Training Strategy: The IWSCL framework is a crucial component of a comprehensive training strategy that also includes PPLG and an instance classifier. This tripartite approach synthesizes the strengths of each component, enhancing the model’s robustness and effectiveness for both instance and bag-level classification tasks. The interaction between IWSCL, PPLG, and the instance classifier, underpinned by weak supervision and enhanced through contrastive learning, significantly improves the training process. It effectively utilizes pseudo-labels to overcome their inherent noise issues and enhance classification performance.
VI-D Novelty of the Proposed PPLG:
Innovation in WSI Classification: WSI classification introduces distinct challenges, notably the reliance on bag-level labels due to the absence of instance-level annotations in a weakly supervised setting. Our PPLG approach overcomes this by generating high-quality pseudo-labels for instance-level training, crucial for accurate predictions at both the instance and bag levels. This represents the first integration of instance-level weakly supervised contrastive learning with prototype learning and an instance classifier in a cohesive training strategy, effectively using true negative instances from negative bags to guide learning.
Innovation Among Related Works: Existing research, as outlined in Section II-D, utilizes prototype methods that typically extract patch features, develop prototypes through various clustering or similarity comparisons, and aggregate them using attention scores. However, these approaches often focus primarily on enhancing bag features and may lack precision in detailed instance classification. They also tend to rely heavily on attention scores, which can inaccurately identify challenging positive instances. Additionally, the integration of feature-level contrastive learning is frequently missing, impacting overall performance. In contrast, our PPLG strategy integrates instance contrastive learning and prototype learning with an instance classifier. This method effectively utilizes classifier guidance and incorporates true negative instances, significantly enhancing pseudo-label quality.
Important Role in the Joint Training Strategy: The PPLG is integral to our joint training strategy, which includes IWSCL and the instance classifier. This holistic approach ensures synergistic contributions from each component, enhancing the overall model efficacy with the PPLG strategy playing a crucial role in refining the pseudo-labels for training the instance classifier.
VII Conclusion